# Optimal Procurement Design: A Reduced Form Approach

### v1

@@
A Reduced Form Approach
@@
I am indebted to Andreas Kleiner and Alejandro Manelli for their encouragement and guidance. I wish to thank Nageeb Ali, Hector Chade, Makoto Hanazono, Ilia Krasikov, Claudio Mezzetti, Benny Moldovanu, Juan Ortner, Edward Schlee, Ji Shen, Kunpeng Sun, Alexander Teytelboym, Mark Whitmeyer, Alexander Wolitzky, Renkun Yang, Mengxi Zhang, and Yimeng Zhang for their advice. I am also grateful to the seminar audiences at ASU, CUFE, and Xiamen University for their feedback. All errors are my own.
(April 21, 2025)
Standard procurement models assume that the buyer knows the quality of the good at the time of procurement; however, in many settings, the quality is learned only long after the transaction. We study procurement problems in which the buyer’s valuation of the supplied good depends directly on its quality, which is unverifiable and unobservable to the buyer. For a broad class of procurement problems, we identify procurement mechanisms maximizing any weighted average of the buyer’s expected payoff and social surplus. The optimal mechanism can be implemented by an auction that restricts sellers to submit bids within specific intervals.
@@ 1 Introduction
Procurement is essential for many different types of organizations. For example, hospitals rely on it to acquire medicines and medical devices (Bonaccorsi, Lyon, Pammolli, and Turchetti, 2000), and governments use it to obtain weapons and public services (Lalive, Schmutzler, and Zulehner, 2015). More generally, firms obtain most of their production inputs as well as many components of their product lines from procurement.
@@
In many procurement contexts, the buyer’s valuation of the supplied good depends directly on its quality. However, at the time the procurement contract is signed, the quality is often (at least partially) unverifiable in court and unobservable by the buyer. For example, this issue arises in both medical and government procurement, when acquiring medicines, medical devices, new weapons, or transportation systems. Although the buyer may have an estimate of the good’s quality before signing the contract, the true quality often remains uncertain until long after the contract is executed. As pointed out by Manelli and Vincent (1995) and Lopomo et al. (2023), in such contexts, buyers who run standard procurement auctions may harbor “quality concerns”—that is, aggressively bidding sellers could be offering low-quality goods, potentially leading to undesirable outcomes.
@@
In this paper, we ask the following question: how to optimally design the procurement mechanism when quality concerns are present? Under mild regularity conditions, a bid-restricted auction (BRA) turns out to be optimal for the buyer. A BRA is similar to a second-price auction, which induces an equilibrium in weakly dominant strategies. The primary difference is that sellers are restricted to bidding within a collection of intervals; see Figure 1 for an illustration. The seller with the lowest bid wins the auction and supplies the good, with ties broken uniformly at random. In most cases, the winning seller receives a payment equal to the second lowest bid; however, in special circumstances, a “payment reduction rule” is applied, resulting in a payment lower than the second lowest bid. This special rule helps eliminate sellers’ incentives to take advantage of the “gaps” between the bid intervals.
More generally, we consider the problem of maximizing a weighted average of the buyer’s expected payoff and the social surplus, subject to the requirement that the buyer’s expected payoff is nonnegative. This constraint is particularly relevant in government procurement design, as the government typically aims to avoid running a deficit. In this more general setting, an optimal procurement mechanism is still a BRA, although a stochastic reserve price might need to be introduced.
@@
Figure 1: The illustration of a BRA. Sellers who wish to participate in the auction must submit a bid $b$ within the two bid intervals, $[\underline{b}_{1},\overline{b}_{1}]$ and $[\underline{b}_{2},\overline{b}_{2}]$. In other words, bids strictly between $\overline{b}_{1}$ and $\underline{b}_{2}$ are not allowed.
@@
The key insight is that the optimal BRA balances the tension between inducing price competition and alleviating quality concerns. Leveraging the structure of a second-price auction, a BRA fosters price competition within bid intervals, where quality concerns are mild. By prohibiting bids in the “gaps” between bid intervals—where quality concerns are severe—harmful price competition is eliminated. This reduces the likelihood of selecting a seller offering a low-quality good, thereby mitigating quality concerns. Put differently, while “global” quality concerns might necessitate abandoning competitive bidding altogether, “local” quality concerns can effectively be addressed by excluding certain bids from the auction.
@@
We establish the optimality of BRA using a “reduced-form approach.” Specifically, we transform the procurement design problem into choosing an interim allocation rule, also known as a “reduced form” in the literature, that maximizes a linear functional identified by some virtual surplus. An interim allocation rule specifies the expected probability that the buyer procures from each seller based on their reports of their private information. The interim allocation rule must satisfy a monotonicity constraint to ensure truthtelling and Border’s (1991) condition to guarantee feasibility. We then apply techniques from linear optimization under a majorization constraint, developed by Kleiner, Moldovanu, and Strack (2021), to solve for the optimal interim allocation. We complete the argument by verifying that a BRA induces the optimal interim allocation.
@@
The study of procurement problems with unverifiable and unobservable qualities is pioneered by Manelli and Vincent (1995). They point out that in many cases, a standard procurement auction may perform poorly due to quality concerns; instead, it may be optimal to sequentially render take-it-or-leave-it offers to potential sellers. Manelli and Vincent (2004) show that some “hybrid mechanisms,” which combine elements of sequential offers and auctions, can be optimal in specific procurement settings. In contrast to these studies, our work identifies procurement mechanisms that maximize any weighted average of the buyer’s expected payoff and social surplus across a broad class of procurement problems.
@@
Lopomo, Persico, and Villa (2023) show that when the buyer’s virtual surplus is single-peaked, a mechanism called the Lowball Lottery Auction (LoLA) is optimal. LoLA differs from a standard second-price procurement auction with a reserve price only in that sellers are not allowed to bid below a certain “floor price.” Our work generalizes Lopomo et al. (2023) by identifying the optimal procurement mechanisms in a broader class of environments; [^1] This is not merely a technical curiosity; violations of their assumption may naturally arise due to the nature of the procurement setting (see Example 1 for a concrete example). indeed, when the buyer’s virtual surplus is single-peaked, the optimal BRA is a LoLA. More importantly, our results elucidate that what matters is not merely whether to use a floor price, but rather to disallow sellers from submitting bids in certain intervals where quality concerns are severe. Furthermore, Lopomo et al. (2023) do not impose the restriction that the buyer’s expected payoff must be nonnegative, and therefore do not offer insights into how this constraint should be addressed, an issue that can be relevant in practice. [^2] Interested readers are directed to Che (2008) for a survey of the procurement literature and Lopomo et al. (2023) for a review of recent developments on procurement design under quality concerns.
@@
The remainder of this paper is organized as follows. Section 2 sets up and transforms the optimal procurement design problem. For clearer insights, in Section 3 we study the special case of maximizing the buyer’s expected payoff; Section 4 tackles the general problem of maximizing any weighted average of the buyer’s payoff and the social surplus, with the constraint that the buyer’s payoff has to be bounded below by zero. Section 5 discusses an extension of the main model as well as a dynamic implementation of the optimal mechanism. Section 6 concludes.
@@ 2 The procurement problem
The model, which is essentially the same as in Manelli and Vincent (1995, 2004), consists of one buyer and $n>1$ symmetric potential sellers. The buyer would like to procure one unit of a product from one of the potential sellers. Each seller $s$ has private information $q_{s}\in[0,1]$. We will refer to $q_{s}$ as the quality of the product offered by seller $s$; it can also be interpreted as seller $s$’s cost or reservation value. [^3] These interpretations are further elaborated in Section 5.1. Qualities are independently and identically distributed according to a continuous density function $f$; we denote the corresponding cumulative distribution function by $F$. We also assume that $f(q)>0$ for $q\in(0,1]$, and $f(0)=0$ only if $\lim_{q\to 0}\left(F(q)/f(q)\right)=0.$ Since we assume that the potential sellers are symmetric, we often suppress the subscript of a seller’s quality.
All agents in our model are expected utility maximizers. If the buyer procures the good from a seller, and a transfer $t$ is made, the seller’s payoff is $t-q$. If a seller does not sell, her payoff is zero. The buyer’s valuation for a good of quality $q$ is a continuous function $v(q)$; we assume that $v(0)\geq 0$. If the buyer makes a transfer $t$ and receives an object of quality $q$, her payoff is $v(q)-t$; if no trade occurs, the buyer’s payoff is zero.
By the revelation principle, it suffices to focus on direct mechanisms. A direct mechanism is characterized by a pair of functions $p_{s}:[0,1]^{n}\to[0,1]$ and $t_{s}:[0,1]^{n}\to\mathbb{R}$ for each seller $s$. If the sellers report $\bm{q}:=(q_{1},\ldots,q_{n})$, the buyer procures from seller $s$ with probability $p_{s}(\bm{q})$, and she makes transfer $t_{s}(\bm{q})$ to seller $s$. Because the buyer wishes to acquire (at most) one unit of the product, for each $\bm{q}\in[0,1]^{n}$, the feasibility constraint must hold:
$\sum_{s=1}^{n}p_{s}(\bm{q})\leq 1;$ (F)
(F) requires that the probability that the buyer buys from one of the potential sellers is less than or equal to 1.
@@
$P_{s}(q^{\prime}_{s}):=\int p_{s}\left(q^{\prime}_{s},\bm{q}_{-s}\right)f^{n-1}\left(\bm{q}_{-s}\right)\,\mathrm{d}\bm{q}_{-s},$
where $\bm{q}_{-s}:=(q_{1},\ldots,q_{s-1},q_{s+1},\ldots,q_{n})$, and $f^{n-1}(\bm{q}_{-s}):=\prod_{k\neq s}f(q_{k})$; she would expect to receive a monetary transfer of
$T_{s}(q^{\prime}_{s}):=\int t_{s}\left(q^{\prime}_{s},\bm{q}_{-s}\right)f^{n-1}\left(\bm{q}_{-s}\right)\,\mathrm{d}\bm{q}_{-s}.$
@@
$\pi_{s}(q^{\prime}_{s}\mid q_{s}):=T_{s}(q^{\prime}_{s})-q_{s}P_{s}(q^{\prime}_{s});$
and we let $\pi_{s}(q_{s}):=\pi_{s}(q_{s}\,|\,q_{s})$. We say that a direct mechanism $\{p_{s},t_{s}\}_{s=1}^{n}$ is incentive compatible if for each seller $s$, all $q^{\prime}_{s}\in[0,1]$, and (almost) all $q_{s}\in[0,1]$, $\pi_{s}(q_{s})\geq\pi_{s}(q^{\prime}_{s}\mid q_{s})$; say that it is individually rational for the sellers if $\pi_{s}(q_{s})\geq 0$ for each seller $s$ and $q_{s}\in[0,1]$. Finally, the buyer’s expected payoff under direct mechanism $\{p_{s},t_{s}\}_{s=1}^{n}$ is
$\pi_{b}:=$ $\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)p_{s}(\bm{q})-t_{s}(\bm{q})\right]f^{n}(\bm{q})\,\mathrm{d}\bm{q},$ (1)
where $f^{n}(\bm{q}):=\prod_{s=1}^{n}f(q_{s})$.
Lemma 1 characterizes the set of incentive compatible direct mechanisms and also eliminates transfers from the buyer’s expected payoff. The proof is standard and hence omitted.
@@
Let $\{p_{s}(\cdot)\}_{s=1}^{n}$ be a collection of allocation functions, where $p_{s}:[0,1]^{n}\to[0,1]$ satisfying (F).
@@
    There exists a collection of transfers $\{t_{s}(\cdot)\}_{s=1}^{n}$ such that $\{(p_{s}(\cdot),t_{s}(\cdot)\}_{s=1}^{n}$ is incentive compatible if and only if for each $s=1,\ldots,n$, $P_{s}(\cdot)$ is decreasing.
@@
    For any incentive compatible direct mechanism $\{(p_{s}(\cdot),t_{s}(\cdot)\}_{s=1}^{n}$,
@@
      it is individually rational for the sellers if and only if $\pi_{s}(1)\geq 0$; and
@@
      $\pi_{b}=\sum_{s=1}^{n}\int_{0}^{1}\left[v\left(q_{s}\right)-q_{s}-\frac{F\left(q_{s}\right)}{f\left(q_{s}\right)}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}-\sum_{s=1}^{n}\pi_{s}(1).$
Without loss of generality, we can set $\pi_{s}(1)=0$ for all $s=1,\ldots,n$. Then by Lemma 1, the weighted average of the buyer’s expected payoff (with weight $\gamma$) and the social surplus (with weight $1-\gamma$) can be written as
$\gamma\sum_{s=1}^{n}\int_{0}^{1}\left[v\left(q_{s}\right)-q_{s}-\frac{F\left(q_{s}\right)}{f\left(q_{s}\right)}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}+(1-\gamma)\sum_{s=1}^{n}\int_{0}^{1}\left[v\left(q_{s}\right)-q_{s}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}$
$=\sum_{s=1}^{n}\int_{0}^{1}\left[v\left(q_{s}\right)-q_{s}-\gamma\frac{F(q_{s})}{f(q_{s})}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}$ (2)
Therefore, we consider the following maximization problem:
$\max_{\{p_{s}\}_{s=1}^{n}}~{}~{}~{}$ $~{}~{}\sum_{s=1}^{n}\int_{0}^{1}\left[v\left(q_{s}\right)-q_{s}-\gamma\frac{F(q_{s})}{f(q_{s})}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}$
[…]
@@
$~{}~{}\sum_{s=1}^{n}\int_{0}^{1}\left[v\left(q_{s}\right)-q_{s}-\frac{F\left(q_{s}\right)}{f\left(q_{s}\right)}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}\geq 0.$
The inequality constraint requires that the buyer’s expected payoff must be nonnegative.
@@
$\max_{\{p_{s}\}_{s=1}^{n}}~{}~{}~{}$ $~{}~{}\sum_{s=1}^{n}\int_{0}^{1}\left[v\left(q_{s}\right)-q_{s}-\frac{F\left(q_{s}\right)}{f\left(q_{s}\right)}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}$
[…]
@@
The inequality constraint is not needed because setting $p_{s}(\bm{q})=0$ for all $\bm{q}\in[0,1]^{n}$ and for $s=1,\ldots,n$ is feasible, ensuring that the value of the problem is nonnegative.
2.1 Detour: Majorization
Let $f,g\in L^{1}(0,1)$ be decreasing. Say that $f$ majorizes $g$, denoted by $g\prec f$, if the following two conditions hold:
$\int_{0}^{x}g(s)\,\mathrm{d}s$ $\leq\int_{0}^{x}f(s)\,\mathrm{d}s~{}~{}\text{ for all }x\in[0,1],$ (3)
$\int_{0}^{1}g(s)\,\mathrm{d}s$ $=\int_{0}^{1}f(s)\,\mathrm{d}s.$ (4)
Say that $f$ weakly majorizes $g$, denoted by $g\prec_{w}f$, if (3) holds (but not necessarily (4)). In what follows, we make use of some results on maximizing a linear functional under a majorization constraint developed by Kleiner et al. (2021). For readers’ convenience, we include these results in Appendix A.
2.2 Transforming the problem: A reduced form approach
One novelty of this paper is that, instead of using a duality approach to characterize the optimal ex post allocation rule, we adopt a reduced form approach: we solve for the optimal interim allocation and identify a trading mechanism that implements it.
To this end, note that the objective function (2) can be written as, in terms of interim allocation probabilities,

### v2

# Optimal Procurement Design: A Reduced-Form Approach

@@
A Reduced-Form Approach
@@
I am indebted to Andreas Kleiner and Alejandro Manelli for their invaluable encouragement and guidance, and I wish to thank the coeditor, Alexander Wolitzky, and four anonymous referees for their excellent editorial comments. For helpful feedback and advice, I am grateful to Nageeb Ali, Hector Chade, Makoto Hanazono, Ravi Jagadeesan, Yixing Li, Simon Loertscher, Dong-Hyuk Kim, Ilia Krasikov, Benny Moldovanu, Juan Ortner, Edward Schlee, Ji Shen, Kunpeng Sun, Alexander Teytelboym, Mark Whitmeyer, Renkun Yang, Mengxi Zhang, and Yimeng Zhang, with special thanks to Claudio Mezzetti. I also thank the seminar audiences at ASU, Auburn, CUFE, UC Riverside, and Xiamen University, as well as the participants at AETW (Auckland). All errors are my own.
(March 8, 2026)
Standard procurement models assume that the buyer knows the quality of the good at the time of procurement; however, in many settings, the quality is learned only long after the transaction. We study procurement problems in which the buyer’s valuation of the supplied good depends directly on its quality, which is unverifiable and unobservable to the buyer. For a broad class of procurement problems, we identify procurement mechanisms maximizing any weighted average of the buyer’s expected payoff and social surplus. The optimal mechanism can be implemented via an auction that restricts sellers to submitting bids within specific intervals.
@@ 1 Introduction
Procurement is essential for many different types of organizations. Unlike typical market purchases, in which goods are often standardized and readily available from many sellers, procurement typically involves acquiring custom goods or services from a limited number of potential suppliers. For example, hospitals rely on procurement to acquire medicines and medical devices (blpt), governments use it to obtain weapons and public services (lsz), and firms obtain most of their production inputs as well as many components of their product lines via procurement.
@@
In such procurement contexts, the buyer’s valuation of the supplied good or service depends directly on its quality. However, at the time the procurement contract is signed, the quality is often (at least partially) unverifiable in court and unobservable by the buyer. Indeed, although the buyer may have an initial estimate of quality, the true stability of a custom software system or the true reliability of a transportation system, for example, often remains uncertain until long after the contract is executed. As pointed out by mv95 and lpv, in such contexts, buyers who run standard procurement auctions may harbor “quality concerns”—that is, concerns that aggressively bidding sellers could be offering low-quality goods, potentially leading to undesirable outcomes.
@@
In this paper, we ask the following question: What is the optimal design of the procurement mechanism when quality concerns are present? Under mild regularity conditions, a bid-restricted auction (BRA) turns out to be optimal for the buyer. Like a second-price auction, a BRA induces an equilibrium in weakly dominant strategies. The primary difference is that in a BRA sellers are restricted to bidding within a collection of intervals; see Figure 1 for an illustration. The seller with the lowest bid wins the auction and supplies the good, and ties are broken uniformly at random. Because bids are only permitted within these intervals, making a bid that lies in a “gap” between intervals is infeasible. Accordingly, a seller who would like to submit such a bid “rounds up” to the smallest permissible bid above the gap (i.e., the lower endpoint of the next interval) so that the entire excluded range of bids collapses into a single bid. Whenever this bid is the lowest bid submitted, the tie-breaking rule applies, yielding a natural form of pooling (random allocation) over the gap. In most cases, the winning seller receives a payment equal to the second-lowest bid; however, in special circumstances, a “payment reduction rule” is applied, resulting in a payment lower than the second-lowest bid. This special rule helps eliminate incentives to bid into the adjacent lower interval to break a tie and thereby increase the probability of winning; as a result, the bid restrictions sustain the intended pooling over the gaps.
More generally, we consider the problem of maximizing any weighted average of the buyer’s expected payoff and the social surplus, subject to the requirement that the buyer’s expected payoff is non-negative. This requirement can be interpreted as the buyer’s individual rationality constraint: the buyer can always decline to procure when her expected net value is negative. It can also be viewed as a reduced-form “value-for-money” screen used in procurement appraisal. [^1] Although quality-related benefits are not literal budget items, many public-sector procurement appraisal frameworks explicitly allow for non-financial performance considerations and other “unmonetized” effects to enter value-for-money comparisons; see, for example, Australian Government, Commonwealth Procurement Rules 2025; HM Treasury, Green Book Supplementary Guidance: Value for Money; and U.S. OMB, Circular A-94. In this more general setting, an optimal procurement mechanism is still a BRA, although a random reserve price might need to be introduced.
@@
Figure 1: Illustration of a BRA. Sellers who wish to participate in the auction must submit a bid $b$ within the two bid intervals, $\underline{b}_{1},\overline{b}_{1}]$ and $\underline{b}_{2},\overline{b}_{2}]$. In other words, bids strictly between $\overline{b}_{1}$ and $\underline{b}_{2}$ are not allowed.
@@
The key insight of this result is that the optimal BRA balances the tension between inducing price competition and alleviating quality concerns. It does so by fostering price competition within bid intervals, where quality concerns are mild; it simultaneously prohibits bids in the gaps between intervals, with the aim of dampening incentives for aggressive price competition in the range in which such competition is most problematic from a quality perspective. Restricting bids to suitably chosen intervals reconciles quality concerns with the simplicity of a second-price-like auction format. Put differently, while “global” quality concerns might necessitate abandoning competitive bidding altogether, “local” quality concerns can effectively be addressed by excluding only those bids that would otherwise fuel undesirable price competition.
@@
We establish the optimality of BRA using a reduced-form approach. Specifically, we transform the procurement design problem into the problem of choosing an interim allocation rule, also known as a “reduced form” in the literature, that maximizes a linear functional identified by some virtual surplus. An interim allocation rule specifies the expected probability that the buyer procures a good from a particular seller based on her report of her private information. The interim allocation rule must satisfy a monotonicity constraint to ensure truth-telling and border1991’s (border1991) condition to guarantee feasibility. We then apply techniques from linear optimization under a majorization constraint, developed by kms, to solve for the optimal interim allocation. We complete the argument by verifying that a BRA induces the optimal interim allocation.
@@
The study of procurement problems with unverifiable and unobservable quality was pioneered by mv95.222See che2008 for a broader survey of the procurement literature; for a summary of recent papers on the performance of certain ad hoc mechanisms in the presence of quality concerns, see lpv. They point out that in many cases, a standard procurement auction may perform poorly due to quality concerns; instead, it may be optimal to sequentially render take-it-or-leave-it offers to potential sellers. mv04 show that some “hybrid mechanisms,” which combine elements of sequential offers and auctions, can be optimal in specific procurement settings. In contrast to these studies, our work identifies procurement mechanisms that maximize any weighted average of the buyer’s expected payoff and social surplus across a broad class of procurement problems.
@@
lpv show that when the buyer’s virtual surplus is single-peaked, a mechanism called the lowball lottery auction (LoLA) is optimal. LoLA differs from a standard second-price procurement auction with a reserve price only in that sellers are not allowed to bid below a certain “floor price.” Our work generalizes lpv by identifying the optimal procurement mechanisms in a broader class of environments; [^3] This is not merely a technical curiosity; violations of their assumption may naturally arise due to the nature of the procurement setting (see Example 1 for a concrete example). indeed, when the buyer’s virtual surplus is single-peaked, the optimal BRA is a LoLA. More importantly, our results reveal that what matters is not merely whether a floor price is used but whether sellers are disallowed from submitting bids in certain intervals in which quality concerns are severe. Furthermore, lpv do not impose the restriction that the buyer’s expected payoff must be non-negative and therefore do not offer insights into how this constraint should be addressed, an issue that can be relevant in practice.
In a BRA, the gaps between bid intervals can be interpreted as “pooling regions.” Other papers in the procurement literature also feature pooling regions as a part of their optimal mechanisms. burguet2012limited and chillemi2014optimal study optimal procurement mechanisms under the risk of default or contract breach. They show, among other findings, that pooling can be useful in reducing the probability of default. cck study procurement design when sellers may collude without using side payments. They show that a collusion-proof mechanism requires pooling on intervals on which the concavified type distribution is affine. [^4] In the context of college admission contests, krishna2026pareto determine the intervals on which pooling introduces Pareto improvement relative to separation by concavifying the type distribution. In this paper, by contrast, pooling regions are introduced to mitigate quality concerns, and they are identified by concavifying certain virtual surplus functions.
@@
The remainder of this paper is organized as follows. Section 2 sets up and transforms the optimal procurement design problem. In Section 3 we study the special case of maximizing the buyer’s expected payoff; Section 4 discusses the general problem of maximizing any weighted average of the buyer’s payoff and the social surplus, with the constraint that the buyer’s payoff is bounded below by zero. Section 5 concludes.
@@ 2 The procurement problem
The model, which is essentially the same as that in mv95; mv04, consists of one buyer and $n>1$ symmetric potential sellers. The buyer would like to procure one unit of a product from one of the potential sellers. Each seller $s$ has private information, $q_{s}\in[0,1]$. We refer to $q_{s}$ as the quality of the product offered by seller $s$; it can also be interpreted as seller $s$’s cost or reservation value. [^5] These interpretations are discussed further in Appendix D. Qualities are independently and identically distributed according to a continuous density function $f$; we denote the corresponding cumulative distribution function by $F$. We also assume that $f(q)>0$ for $q\in(0,1]$ and that $f(0)=0$ only if $\lim_{q\to 0}\left(F(q)/f(q)\right)=0.$ Since we assume that the potential sellers are symmetric, we often suppress the subscript of a seller’s quality.
All agents in our model are expected utility maximizers. If the buyer procures a good from a seller and a transfer $t$ is made, the seller’s payoff is tqt-q. If a seller does not sell, her payoff is zero. The buyer’s valuation for a good of quality $q$ is a continuous function $v(q)$; we assume that $v(0)\geq 0$. If the buyer makes a transfer $t$ and receives an object of quality $q$, her payoff is v(q)tv(q)-t; if no trade occurs, the buyer’s payoff is zero.
By the revelation principle, it suffices to focus on direct mechanisms. A direct mechanism is characterized by a pair of functions, $p_{s}:[0,1]^{n}\to[0,1]$ and $t_{s}:[0,1]^{n}\to\mathbb{R}$, for each seller $s$. If the sellers report $\bm{q}\coloneqq(q_{1},\ldots,q_{n})$, the buyer procures from seller $s$ with probability $p_{s}(\bm{q})$ and she makes transfer $t_{s}(\bm{q})$ to seller $s$; we call $p_{s}(\cdot)$ the ex-post allocation probability for seller $s$. Because the buyer wishes to acquire (at most) one unit of the product, for each $\bm{q}\in[0,1]^{n}$, the following feasibility constraint must hold:
$\sum_{s=1}^{n}p_{s}(\bm{q})\leq 1.$ (F)
Condition (F) requires that the probability that the buyer buys from one of the potential sellers is less than or equal to 1.
@@
$P_{s}(q^{\prime}_{s})\coloneqq\int p_{s}\left(q^{\prime}_{s},\bm{q}_{-s}\right)f^{n-1}\left(\bm{q}_{-s}\right)\,\mathrm{d}\bm{q}_{-s},$
where $\bm{q}_{-s}\coloneqq(q_{1},\ldots,q_{s-1},q_{s+1},\ldots,q_{n})$ and $f^{n-1}(\bm{q}_{-s})\coloneqq\prod_{k\neq s}f(q_{k})$; she would expect to receive a monetary transfer of
$T_{s}(q^{\prime}_{s})\coloneqq\int t_{s}\left(q^{\prime}_{s},\bm{q}_{-s}\right)f^{n-1}\left(\bm{q}_{-s}\right)\,\mathrm{d}\bm{q}_{-s}.$
@@
$\pi_{s}(q^{\prime}_{s}\mid q_{s})\coloneqq T_{s}(q^{\prime}_{s})-q_{s}P_{s}(q^{\prime}_{s}),$
and we let $\pi_{s}(q_{s})\coloneqq\pi_{s}(q_{s}\,|\,q_{s})$. For a given direct mechanism $\{p_{s},t_{s}\}_{s=1}^{n}$ , the (expected) social surplus is
$\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)-q_{s}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q},$
where $f^{n}(\bm{q})\coloneqq\prod_{s=1}^{n}f(q_{s})$; the buyer’s expected payoff is
$\pi_{b}\coloneqq$ $\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)p_{s}(\bm{q})-t_{s}(\bm{q})\right]f^{n}(\bm{q})\,\mathrm{d}\bm{q}.$ (1)
We say that a direct mechanism $\{p_{s},t_{s}\}_{s=1}^{n}$ is incentive-compatible if every seller truthfully reports her quality; formally, for each seller $s$, all $q^{\prime}_{s}\in[0,1]$ and (almost) all $q_{s}\in[0,1]$, $\pi_{s}(q_{s})\geq\pi_{s}(q^{\prime}_{s}\mid q_{s})$. We say that it is individually rational if the buyer and all sellers are willing to participate in the mechanism; that is, $\pi_{b}\geq 0$ and $\pi_{s}(q_{s})\geq 0$ for each seller $s$ and $q_{s}\in[0,1]$.
Lemma 1 characterizes the set of incentive-compatible direct mechanisms and eliminates transfers from the buyer’s expected payoff. The proof is standard and hence omitted.
@@
Let $\{p_{s}(\cdot)\}_{s=1}^{n}$ be a collection of ex-post allocation probabilities, where $p_{s}:[0,1]^{n}\to[0,1]$, satisfying (F).
@@
    There exists a collection of transfers $\{t_{s}(\cdot)\}_{s=1}^{n}$ such that $\{(p_{s}(\cdot),t_{s}(\cdot))\}_{s=1}^{n}$ is incentive-compatible if and only if for each $s=1,\ldots,n$, $P_{s}(\cdot)$ is decreasing. [^6] We use “decreasing” in the weak sense: “strict” will be added whenever needed.
@@
    For any incentive-compatible direct mechanism $\{(p_{s}(\cdot),t_{s}(\cdot))\}_{s=1}^{n}$,
@@
      it is individually rational if and only if $\pi_{b}\geq 0$ and $\pi_{s}(1)\geq 0$ for each $s=1,\ldots,n$; and
@@
      $\pi_{b}=\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)-q_{s}-\frac{F\left(q_{s}\right)}{f\left(q_{s}\right)}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}-\sum_{s=1}^{n}\pi_{s}(1).$
By Lemma 1, the weighted average of the buyer’s expected payoff (with weight $\gamma$) and the social surplus (with weight $1-\gamma$) can be written as
$\gamma\left\{\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)-q_{s}-\frac{F\left(q_{s}\right)}{f\left(q_{s}\right)}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}-\sum_{s=1}^{n}\pi_{s}(1)\right\}+$
$(1-\gamma)\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)-q_{s}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}$
$=\,$ $\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)-q_{s}-\gamma\frac{F(q_{s})}{f(q_{s})}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}-\gamma\sum_{s=1}^{n}\pi_{s}(1).$
Since individual rationality for the sellers is equivalent to $\pi_{s}(1)\geq 0$ for each $s=1,\ldots,n$, to maximize this weighted average, we set $\pi_{s}(1)=0$ for all $s=1,\ldots,n$. [^7] For $\gamma>0$, the condition $\pi_{s}(1)=0$ for all $s=1,\ldots,n$ is necessary for a mechanism to be optimal. For $\gamma=0$, setting $\pi_{s}(1)=0$ does not change the social surplus but does relax the buyer’s individual rationality constraint. Therefore, we consider the following maximization problem:
$\max_{\{p_{s}\}_{s=1}^{n}}$~~~ $~~\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)-q_{s}-\gamma\frac{F(q_{s})}{f(q_{s})}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}$ (2)
subject to (F)
@@
$~~\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)-q_{s}-\frac{F\left(q_{s}\right)}{f\left(q_{s}\right)}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}\geq 0.$
The inequality constraint requires that the buyer’s expected payoff is non-negative.
@@
$\max_{\{p_{s}\}_{s=1}^{n}}$~~~ $~~\sum_{s=1}^{n}\int_{[0,1]^{n}}\left[v\left(q_{s}\right)-q_{s}-\frac{F\left(q_{s}\right)}{f\left(q_{s}\right)}\right]p_{s}(\bm{q})f^{n}(\bm{q})\,\mathrm{d}\bm{q}$
subject to (F)
@@
The inequality constraint is not needed because setting $p_{s}(\bm{q})=0$ for all $\bm{q}\in[0,1]^{n}$ and for $s=1,\ldots,n$ is feasible, ensuring that the value of the problem is non-negative.
2.1 Transforming the problem: A reduced-form approach
Instead of solving directly for optimal ex-post allocation probabilities (e.g., myerson1981; mv95), we take a reduced-form approach. We first solve for the optimal interim allocation probabilities $\{P_{s}(\cdot)\}_{s=1}^{n}$ and then identify a trading mechanism that is consistent with them. This approach is valid because a seller’s report affects her incentives only through the interim allocation probability it generates, and the feasibility constraint can be expressed in terms of interim allocation probabilities as well. Working with interim allocations is also convenient: each $P_{s}(\cdot)$ is a function of a single variable, whereas ex-post allocation probabilities are functions of $n$ variables.
To this end, note that the objective function (2), in terms of interim allocation probabilities, can be written as